Ali Inc. expects to generate free-cash of $200000per year forever. If the firm's cost of capital is 0.15 percent,the firm cost of equity capital is 0.16 the market value of debt is $320000, the market value of preferred stock is $180000, and the company has 120000 shares of stock outstanding. What is the value of Ali's stock?what is the value of the firm

Answer for part 1


what is the value of the c.s?


1
Expert's answer
2021-09-07T07:48:02-0400

"200000 \u00d7 0.15 = 30000"


"0.16 \u00d7 320000 = 51200"


"30000+51200 = 81200"


"= 81200 + 180000 = 261000"


"\frac{261000}{120000}" "= 2.175"

"Value of stock = 2.175"